OpenStudy (anonymous):

find the LCD of, 1over11 , and 51over121 ?

OpenStudy (zzr0ck3r):

is it not 121?

OpenStudy (zzr0ck3r):

this is least common denominator not divisor

OpenStudy (zzr0ck3r):

1/11 + 51/121 = 11/121 + 51/121 = 62/151

OpenStudy (zzr0ck3r):

lcd = 121

OpenStudy (zzr0ck3r):

@FutureMathProfessor you delete everything and it makes me look like im talking to myself.

OpenStudy (zzr0ck3r):

lol